this is from my theory of computation class i need this question is detail ans,

Find a simple and nontrivial characterization of the language {111} ∗{11111} ∗ . Give a proof of the correctness of your answer.
Answer #1

We are given the language {111}*{11111}* which can also be written as {111+11111}* = {1}+17 .

So any string from the language can be written as 1^z where 2 = 3х + 5y and x and y are non-negative integers.

The set {-12 = 3.7 + 5y, x,y EZ and r, y > 0 contains all non-negative integers except 1,2,4,6 and 7.
This is because ged(3,5) = 1 and so we must have integers x and y ​​​satisfying 2= 3.+ 5y, VzEZ . However if we restrict the domain of x and y to non-negative integers we are not getting some integers, and to be precise the integers {1, 2, 4, 6, 7} .

So, we can characterize the language as 1* - {0, 1, 11, 1111, 111111. 1111111 .